# Palm distributions of superposed point processes for statistical inference

### v1

@@ Abstract
Palm distributions play a central role in the study of point processes and their associated summary statistics. In this paper, we characterize the Palm distributions of the superposition of two independent point processes, establishing a simple mixture representation depending on the point processes’ Palm distributions and moment measures. We explore two statistical applications enabled by our main result. First, we consider minimum contrast estimation for superposed point processes based on Ripley’s $K$ function. Second, we focus on the class of shot-noise Cox processes and obtain a tractable expression for the Janossy density which leads to maximum likelihood estimation via a novel expectation-maximization algorithm. Both approaches are validated through numerical simulations. Extensions to the superposition of multiple point processes, and higher-order Palm distributions, are also discussed.
Keywords: Superposition of point processes; minimum contrast estimation; determinantal point process; shot-noise Cox process; noisy observations; expectation maximization
@@ 1 Introduction
Real-world point patterns often combine several structured components, some regular or inhomogeneous, others explicitly clustered, and may also include random noise. Semiconductor wafer defect maps (Borgons2021), disease-case locations in epidemiology (meyer2017surveillance), cellular network base-station layouts (choi2017gridppp), mixed age tree stands in ecology (ngobieng2011forest), and earthquake aftershock sequences (ogata1998space) can all be regarded as superpositions of two (or more) independent point processes. While the superposition operation is trivial at the model level, from an inferential standpoint, superposed point processes are notoriously awkward. Standard tools such as minimum contrast estimation rely on closed-form expressions for second-order summaries, such as the $J$ function, Ripley’s $K$-function, and Besag’s $L$ function. However, those summaries remain unknown for a generic superposition (MoWaBook03; diggle2013statistical). Hence, practitioners need to rely on complex algorithms often developed on a case-by-case basis (e.g., tanaka2014superposed; xu2018hawkes).
@@
In this work we focus on the Palm distributions of the superposition of two independent point processes. Palm distributions (BaBlaKa) are key mathematical objects in the study of point processes, describing the conditional behavior of a process given the location of one or more of its points or atoms. We establish a mixture representation for the Palm distributions of the superposed process, in which the mixture components are sums of the two original processes and their Palm versions, weighted by their respective mean measures. Our analysis easily extends to the superposition of more than two independent processes.
@@
We demonstrate the practical usefulness of our result for statistical inference in two settings. First, we consider fitting a determinantal point process (dpp, Lavancier2015) contaminated by random background noise via minimum contrast estimation. Indeed, through the Palm distributions, it is straightforward to obtain Ripley’s $K$ function, enabling robust and fast inference via minimum contrast. This methodology is particularly relevant in applied contexts such as the analysis of spatial defect structures in semiconductor manufacturing (Borgons2021).
Second, we consider estimating the parameters of the shot noise Cox process (sncp, Mo03Cox), a prominent class of cluster processes. Cluster processes are routinely employed in several applied areas, such as astronomy, materials science, and plant ecology; see, e.g., Illian2008; MoWaBook03. Despite their generality and usefulness, statistical inference for general sncp models is lacking, with most of the contributions focusing on simple sub-classes (see wang2023spatiotemporal, for a recent contribution) or on simulation of the process with fixed parameters (MoWaBook03). An application of our main result yields the higher-order Palm distributions of the sncp, previously unknown in the literature. Building on this result, we obtain an explicit expression for the Janossy density, which in turn enables maximum likelihood estimation via an expectation-maximization algorithm. We demonstrate the effectiveness of such an approach on the gamma sncps with Thomas kernel (MoWaBook03). We also disprove a conjecture by coeurjolly2017tutorial showing that the higher-order Palm distributions of the sncp do not match the law of a sncp.
Although not explored here, our theoretical analysis is helpful in several other statistical contexts. For example, in Bayesian nonparametrics, the superposition of point processes is used to define a prior distribution when data is subdivided into groups (Nip14; griffin2013comparing). Our main result can be leveraged for posterior analysis and numerical computations.
@@ 2.1 Background and notation for point processes
Let $\mathbb{X}$ be a Polish space equipped with the corresponding Borel $\sigma$-algebra $\mathscr{X}$. A point process $\Phi$ on $\mathbb{X}$ can be represented as $\Phi=\sum_{j\geq 1}\delta_{X_{j}}$, where $(X_{j})_{j\geq 1}$ is a sequence of random variables (atoms) taking values in $\mathbb{X}$, and $\delta_{x}$ denotes the Dirac delta mass at $x$. The probability distribution of $\Phi$ is denoted with $\mathbf{P}_{\Phi}$. The number of atoms in $\Phi$ could be either finite or infinite. In the sequel, we will follow the approach of (BaBlaKa), where $\Phi$ is regarded as a random counting measure. See Appendix S1 for further mathematical details, including the $\sigma$-algebra on the space of counting measures.
Let $M_{\Phi}(B)=\mathsf{E}[\Phi(B)]$, for $B\in\mathscr{X}$, be the mean measure of $\Phi$, and define the $k$-th factorial moment measure $M_{\Phi}^{(k)}$ as the mean measure of the $k$-th factorial power of $\Phi$, i.e., of the point process $\Phi^{(k)}$ defined as:
@@
To introduce the notion of Palm distributions, let us define the Campbell measure of $\Phi$, namely $\mathscr{C}_{\Phi}(B\times L):=\mathsf{E}[\Phi(B)\mathbbm{1}(\Phi\in L)]$, where $B\in\mathscr{X}$ and $L$ is an element of the appropriate $\sigma$-algebra on the space of random counting measures (see Appendix S1). Under the assumption that $M_{\Phi}$ is $\sigma$-finite, it can be shown that $\mathscr{C}_{\Phi}$ admits the following representation
$\mathscr{C}_{\Phi}(B\times L)=\int_{B}\mathbf{P}_{\Phi}^{x}(L)M_{\Phi}(\mathrm{d}x),$
where $\{\mathbf{P}_{\Phi}^{x}\}_{x\in\mathbb{X}}$ is the almost surely (a.s.) unique disintegration probability kernel of $\mathscr{C}_{\Phi}$ with respect to $M_{\Phi}$, and it is referred to as the Palm kernel of $\Phi$. For fixed $x\in\mathbb{X}$, $\mathbf{P}_{\Phi}^{x}$ is a probability distribution over the space of counting measures, termed the Palm distribution of $\Phi$ at $x$, and thus it can be identified with the law of a point process $\Phi_{x}\sim\mathbf{P}_{\Phi}^{x}$, which is consequently called a Palm version of $\Phi$ at $x$. By Proposition 3.1.12 in (BaBlaKa), for $M_{\Phi}$-almost all $x\in\mathbb{X}$, the point process $\Phi_{x}$ contains the atom $x$ with probability $1$. This justifies the interpretation of the Palm distribution of $\Phi$ at $x$ as the law of $\Phi$ conditionally to $\Phi$ having an atom at $x$. In addition, the point process $\Phi^{!}_{x}:=\Phi_{x}-\delta_{x}$ is well-defined, and $\Phi^{!}_{x}$ is called a reduced Palm version of $\Phi$ at $x$. Finally, it is possible to extend the definition of Palm distributions to multiple conditioning points $\bm{x}=(x_{1},\ldots,x_{k})\in\mathbb{X}^{k}$. In this case, the Palm distribution of $\Phi$ at $\bm{x}$ is interpreted as the probability distribution of $\Phi$ conditionally to $\Phi$ having $k$ atoms at locations $x_{1},\ldots,x_{k}$. See Appendix S1.
@@ 2.2 Palm distributions of the superposition of independent processes
Consider independent point processes $\Phi_{i}$, $i=1,\ldots,n$. Their superposition $\Phi$ is defined as $\Phi=\sum_{i=1}^{n}\Phi_{i}$, i.e., the union of all the point patterns. The following theorem states the main theoretical result of the paper, which characterizes the Palm distributions of the superposition of two independent point processes.
@@
Theorem 1 admits an intuitive interpretation about the Palm distributions of the superposed process. Conditioning the superposed process $\Phi=\Phi_{1}+\Phi_{2}$ on having a point at $x$, its distribution depends on whether $x$ originates from $\Phi_{1}$ or from $\Phi_{2}$, thus reflecting the respective contributions of each source. The mixture structure in the Palm version $(\Phi_{1}+\Phi_{2})_{x}$ explicitly accounts for the two mutually exclusive scenarios. The mixing probabilities correspond precisely to the probabilities that point $x$ originates from one or the other process.
Remark 1 (Reduced Palm version of superposition).
Exploiting the relation $\Phi_{ix}\stackrel{{\scriptstyle\mbox{\scriptsize d}}}{{=}}\Phi_{ix}^{!}+\delta_{x}$, for $M_{\Phi_{i}}$-almost all $x$, $i=1,2$, we can replace all the Palm versions appearing in the statement of Theorem 1 with their reduced counterparts.
Theorem 1 can be extended in several directions. In particular, in Appendix S2.2, we characterize the Palm distributions of the superposition of more than two independent point processes, as well as the Palm distributions given multiple conditioning points.
3 Inference for corrupted determinantal process via minimum contrast
3.1 Minimum contrast estimation
Minimum contrast estimimation methods (mce, MoWaBook03) are a class of techniques for fitting parametric point process models to observed point patterns. In several common settings where the likelihood of the point process is intractable, mce is the only viable option to perform inference. Let $\Phi$ be a point process whose distribution $\mathbf{P}_{\Phi}$ is parametrized by a set of parameters $\theta$. Let $s(r;\theta)$, $r\geq 0$, be a functional summary statistics of $\Phi$, for which an analytical expression is available, and let $\hat{s}(r)$ be a nonparametric estimate of $s$ based on the observed data. The mce for $\theta$ is obtained as
$\hat{\theta}=\text{argmin}_{\theta}\int_{r_{l}}^{r_{u}}|\hat{s}(r)^{q}-s(r;\theta)^{q}|^{p}\mathrm{d}r.$ (1)
Common choices consider $r_{l}=0$, $q=1/2$, $p=2$ (diggle2013statistical). Among the various options for $s(r;\theta)$, we focus here on Ripley’s $K$-function (ripley1976second), a canonical choice for stationary point processes; see, e.g., BaddeleyRubakTurner2015 and Illian2008. Assuming that $\Phi$ is stationary with intensity $M_{\Phi}\equiv\rho>0$, the $K$-function of $\Phi$ is defined as
$K_{\Phi}(r)=\frac{1}{\rho}\mathsf{E}[\Phi^{!}_{o}(B(o,r))],\quad\text{for }r\geq 0,$

### v2

@@ Abstract
Palm distributions play a central role in the study of point processes and their associated summary statistics. In this paper, we characterize the Palm distributions of the superposition of independent point processes, establishing a simple mixture representation depending on the point processes’ Palm distributions and moment measures. We explore two statistical applications enabled by our main result. First, we consider minimum contrast estimation for corrupted point processes. Second, we investigate the class of shot noise Cox processes and derive explicit expressions for their higher-order Palm distributions. In the finite case, we further obtain a tractable expression for the Janossy density, which plays the role of a likelihood function and thus can be used for new likelihood-based inference strategies. Extensions to the superposition of multiple point processes and to higher-order Palm distributions are also presented.
Keywords: Matérn cluster process, minimum contrast estimation, noisy observations, shot noise Cox process, summary statistics, superposition of point processes
@@ 1 Introduction
Real-world point patterns often combine several structured components, some regular or inhomogeneous, others explicitly clustered, and may also include random noise. Semiconductor wafer defect maps (Borgoni et al., 2021), disease-case locations in epidemiology (Meyer et al., 2017), cellular network base-station layouts (Choi et al., 2017), mixed age tree stands in ecology (Ngo Bieng et al., 2011), and earthquake aftershock sequences (Ogata, 1998) can all be regarded as superpositions of two (or more) independent point processes. While the superposition operation is trivial at the model level, from an inferential standpoint, superposed point processes are notoriously awkward. Standard tools such as minimum contrast estimation rely on closed-form expressions for second-order summaries, such as the $J$-function, Ripley’s $K$-function, and Besag’s $L$-function. However, those summaries remain unknown for a generic superposition (Møller and Waagepetersen, 2003; Diggle, 2013). Hence, practitioners need to rely on complex algorithms often developed on a case-by-case basis (e.g., Tanaka and Ogata, 2014; Xu et al., 2018).
@@
In this work we first focus on the Palm distributions of the superposition of two independent point processes. Palm distributions (Baccelli et al., 2020) are key mathematical objects in the study of point processes, describing the conditional behavior of a process given the location of one or more of its points or atoms. We establish a mixture representation for the Palm distributions of the superposed process, in which the mixture components are sums of the two original processes and their Palm versions, weighted by their respective mean measures. We also show that our analysis easily extends to the superposition of more than two independent processes.
@@
We demonstrate the practical usefulness of our result for statistical inference in two settings. First, we consider fitting a point process contaminated by random background noise via minimum contrast estimation. Indeed, through the Palm distributions, it is straightforward to obtain functional summary statistics, enabling robust and fast inference via minimum contrast. This methodology is particularly relevant in applied contexts such as the analysis of spatial defect structures in semiconductor manufacturing (Borgoni et al., 2021).
Second, we investigate structural and distributional properties of the shot noise Cox process (sncp, Møller, 2003), a prominent class of cluster processes. Cluster processes are routinely employed in several applied areas, including astronomy, materials science, and plant ecology; see, e.g., Illian et al. (2008); Møller and Waagepetersen (2003). Despite their flexibility and wide applicability, several distributional aspects of sncp models remain only partially understood. As an application of our main result, we derive the higher-order Palm distributions of the sncp, which were previously unavailable in the literature. In the case of finite sncps, we further obtain an explicit expression for the Janossy density. Since the Janossy density plays the role of a likelihood function for finite point processes, this result paves the way for new likelihood-based inference strategies in the context of sncp models. We also discuss the use of our results in other statistical contexts, both frequentist and Bayesian.
@@ 2.1 Background and notation for point processes
Let $\mathbb{X}$ be a Polish space equipped with the corresponding Borel $\sigma$-algebra $\mathscr{X}$. A point process $\Phi$ on $\mathbb{X}$ can be represented as $\Phi=\sum_{j\geq 1}\delta_{X_{j}}$, where $(X_{j})_{j\geq 1}$ is a sequence of random variables (atoms) taking values in $\mathbb{X}$, and $\delta_{x}$ denotes the Dirac delta mass at $x$. The probability distribution of $\Phi$ is denoted with $\mathsf{P}_{\Phi}$. The number of atoms in $\Phi$ could be either finite or infinite. In the sequel, we will follow the approach of Baccelli et al. (2020), where $\Phi$ is regarded as a random counting measure. See Section S1 for further mathematical details, including the $\sigma$-algebra on the space of counting measures.
Let $M_{\Phi}(B)=E[\Phi(B)]$, for $B\in\mathscr{X}$, be the mean measure of $\Phi$, and define the $k$-th factorial moment measure $M_{\Phi}^{(k)}$ as the mean measure of the $k$-th factorial power of $\Phi$, i.e., of the point process $\Phi^{(k)}$ defined as:
@@
To introduce the notion of Palm distributions, let us define the Campbell measure of $\Phi$, namely $\mathscr{C}_{\Phi}(B\times L):=E[\Phi(B){1}(\Phi\in L)]$, where $B\in\mathscr{X}$ and $L$ is an element of the appropriate $\sigma$-algebra on the space of random counting measures (see Section S1). Under the assumption that $M_{\Phi}$ is $\sigma$-finite, it can be shown that $\mathscr{C}_{\Phi}$ admits the following representation
$\mathscr{C}_{\Phi}(B\times L)=\int_{B}\mathsf{P}_{\Phi}^{x}(L)M_{\Phi}(\mathrm{d}x),$
where $\{\mathsf{P}_{\Phi}^{x}\}_{x\in\mathbb{X}}$ is the almost surely (a.s.) unique disintegration probability kernel of $\mathscr{C}_{\Phi}$ with respect to $M_{\Phi}$, and it is referred to as the Palm kernel of $\Phi$. For fixed $x\in\mathbb{X}$, $\mathsf{P}_{\Phi}^{x}$ is a probability distribution over the space of counting measures, termed the Palm distribution of $\Phi$ at $x$, and thus it can be identified with the law of a point process $\Phi_{x}\sim\mathsf{P}_{\Phi}^{x}$, which is consequently called a Palm version of $\Phi$ at $x$. By Proposition 3.1.12 in Baccelli et al. (2020), for $M_{\Phi}$-almost all $x\in\mathbb{X}$, the point process $\Phi_{x}$ contains the atom $x$ with probability $1$. This justifies the interpretation of the Palm distribution of $\Phi$ at $x$ as the law of $\Phi$ conditionally to $\Phi$ having an atom at $x$. In addition, the point process $\Phi^{!}_{x}:=\Phi_{x}-\delta_{x}$ is well-defined, and $\Phi^{!}_{x}$ is called a reduced Palm version of $\Phi$ at $x$. Finally, it is possible to extend the definition of Palm distributions to multiple conditioning points $\underline{x}=(x_{1},\ldots,x_{k})\in\mathbb{X}^{k}$. In this case, the Palm distribution of $\Phi$ at $\underline{x}$ is interpreted as the probability distribution of $\Phi$ conditionally to $\Phi$ having $k$ atoms at locations $x_{1},\ldots,x_{k}$. See Section S1.
@@ 2.2 Palm distributions of the superposition of independent processes
From now on, consider independent simple point processes $\Phi_{i}$, $i=1,\ldots,m$. Their superposition $\Phi$ is defined as $\Phi:=\sum_{i=1}^{m}\Phi_{i}$, i.e., the union of all the point patterns. The following theorem states the main theoretical result of the paper, which characterizes the Palm distributions of the superposition of two independent point processes.
@@
The proof is reported in the Appendix at the end of the paper. Theorem 1 admits an intuitive interpretation about the Palm distributions of the superposed process. Conditioning the superposed process $\Phi=\Phi_{1}+\Phi_{2}$ on having a point at $x$, its distribution depends on whether $x$ originates from $\Phi_{1}$ or from $\Phi_{2}$, thus reflecting the respective contributions of each source. The mixture structure in the Palm version $(\Phi_{1}+\Phi_{2})_{x}$ explicitly accounts for the two mutually exclusive scenarios. The mixing probabilities correspond precisely to the probabilities that point $x$ originates from one or the other process. Furthermore, exploiting the relation $\Phi_{ix}\stackrel{{\scriptstyle\mbox{\scriptsize d}}}{{=}}\Phi_{ix}^{!}+\delta_{x}$, for $M_{\Phi_{i}}$-almost all $x$, $i=1,2$, we can replace all the Palm versions appearing in the statement of Theorem˜1 with their reduced counterparts.
To generalize Theorem 1 to $m$ processes and $k$ points, let us introduce latent allocation variables $T_{1},\ldots,T_{k}\in\{1,\ldots,m\}$ with the interpretation that $T_{j}=i$ iff $x_{j}$ was generated from $\Phi_{i}$.
Theorem 2.
Let $\Phi=\sum_{i=1}^{m}\Phi_{i}$ be the superposition of the independent simple point processes $\Phi_{1},\ldots,\Phi_{m}$, and $\underline{x}=(x_{1},\ldots,x_{k})\in\mathbb{X}^{k}$, with pairwise distinct coordinates. Assume that the $k$-th factorial moment measure $M_{\Phi}^{(k)}$ is $\sigma$-finite, and that $M_{\Phi_{i}}^{(r)}$ admits a density $\rho_{\Phi_{i}}^{(r)}$ w.r.t. $\mu^{\otimes r}$, namely the $r$-fold product of a reference measure $\mu$, for all $i=1,\ldots,m$ and $r\geq 1$. For $\underline{T}:=(T_{1},\ldots,T_{k})$, define $\underline{x}_{i}=(x_{j}:T_{j}=i)$ and denote by $n_{i}$ the cardinality of $\underline{x}_{i}$. Then, for $M_{\Phi}^{(k)}$-almost all $\underline{x}$,
$\Phi^{!}_{\underline{x}}\,|\,(\underline{T}=\underline{t})\ \stackrel{{\scriptstyle\mbox{\scriptsize d}}}{{=}}\ \sum_{i=1}^{m}(\Phi_{i})^{!}_{\underline{x}_{i}},\qquad P(\underline{T}=\underline{t})\propto\prod_{i=1}^{m}\rho_{\Phi_{i}}^{(n_{i})}\big(\underline{x}_{i}\big).$
where, conditionally on $\underline{T}=\underline{t}$, the $(\Phi_{i})^{!}_{\underline{x}_{i}}$’s are mutually independent and, when $n_{i}=0$, $(\Phi_{i})^{!}_{\underline{x}_{i}}=\Phi_{i}$ and $\rho_{\Phi_{i}}^{(n_{i})}\equiv 1$.
3 Inference for corrupted processes via minimum contrast
3.1 Summary statistics for superposed point processes
An application of Theorem 1 yields closed-form and interpretable expressions for commonly used summary statistics. Here we concentrate on two summary statistics for stationary point processes, several additional examples are provided in the supplementary materials. Throughout this section, we assume that $\Phi=\Phi_{1}+\Phi_{2}$ where the $\Phi_{j}$’s are stationary with intensity $\rho_{j}$, and let $\rho=\rho_{1}+\rho_{2}$.
Perhaps the most commonly used summary statistic is Ripley’s $K$-function (Ripley, 1976). Thanks to Theorem 1, for a superposed point process $\Phi$ this takes the following form, for $r\geq 0$,
$K_{\Phi}(r):=\frac{1}{\rho}E[\Phi^{!}_{o}(B(o,r))]=\frac{1}{\rho}\left[K_{\Phi_{1}}(r)\frac{\rho_{1}^{2}}{\rho}+K_{\Phi_{2}}(r)\frac{\rho_{2}^{2}}{\rho}+2|B(o,r)|\frac{\rho_{1}\rho_{2}}{\rho}\right],$ (1)